OpenStudy (anonymous):

i don't understand how to do these i need help 9-2/3x=-1 find out x

OpenStudy (anonymous):

helllllp

jhonyy9 (jhonyy9):

x is on denominator or ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes sorry

jhonyy9 (jhonyy9):

27x-2 = -3x 30x=2 x=2/30 x=1/15

jhonyy9 (jhonyy9):

ok ?